# Double Successive Over-Relaxation Q-Learning with an Extension to Deep Reinforcement Learning

### v1

@@ IV-A Over-estimation bias of SOR Q-learning
Bias $=E[w\max_{i}\mu_{i}+(1-w)\max_{i}\mu_{i}]-\max_{i}E[\mu_{i}]$
$=E[\max_{i}\mu_{i}]-\max_{i}E[\mu_{i}]\geq 0.$
Therefore, with probability $p^{i}_{ia}$, the SORQL algorithm overestimates the true value. Similarly, we can see that the MF-SORQL will also suffer from over-estimation bias. To this end, the techniques developed in double Q-learning prove useful. More precisely, in double Q-learning, for a set of random variables $\{X_{i}\}^{d}_{i=1}$, to evaluate $\max_{i}E[X_{i}]$, two sets of unbiased estimators, namely $\{\mu^{A}_{i}\}^{d}_{i=1}$, and $\{\mu^{B}_{i}\}^{d}_{i=1}$ are used. Further, it is shown that for $a^{*}$ such that $\mu^{A}_{a^{*}}=\max_{i}\mu^{A}_{i}$, $E[\mu^{B}_{a^{*}}]\leq\max_{i}E[X_{i}]$ (Lemma 1, [8]). Therefore, Bias $=E[\mu^{B}_{a^{*}}]-\max_{i}E[\mu^{B}_{i}]\leq 0$, consequently, double Q-learning underestimates the true value. Note that underestimation is not a complete solution to the over-estimation problem of Q-learning. However, this modification to Q-learning has proven to be more useful in many situations [24]. Now, for the proposed DSORQL, let $\{\mu^{A}_{i}\}^{d}_{i=1}$, $\{\mu^{B}_{i}\}^{d}_{i=1}$, and $\{\zeta^{A}_{j}\}^{d}_{j=1}$, $\{\zeta^{B}_{j}\}^{d}_{j=1}$ be the unbiased estimators of $\{X_{i}\}^{d}_{i=1}$, and $\{Y_{j}\}^{d}_{j=1}$, respectively. Therefore, the bias is given by
$\text{Bias}=E[w\mu^{B}_{a^{*}}+(1-w)\zeta^{B}_{b^{*}}]-\max_{i}E[\mu^{B}_{i}],$ (9)
@@ V Numerical Experiments
In this section, the tabular and deep RL versions of the proposed algorithms are compared on benchmark examples. The code for all the numerical experiments in this manuscript is publicly available and can be accessed in the author’s GitHub repository [25].
Figure 1: SOR with $w=1.3$.
Figure 2: SOR with $w=5$.
Figure 3: SOR with $w=10$.
Figure 4: SOR with $w=w^{*}=1/(1-\gamma)$.
@@ V-A Tabular version
This subsection provides numerical comparisons between Q-learning (QL), double Q-learning (DQL), model-free SOR Q-learning (MF-SORQL), and model-free double SOR Q-learning (MF-DSORQL). More specifically, we compare the above-mentioned algorithms on benchmark examples used in [8], namely roulette and grid world. Note that the step size for all the algorithms is the same, and all the other parameters used in the experiments are similar to [8].
V-A1 Roulette
In roulette, a player can choose from $170$ different betting actions, such as betting on a specific number or on the color black or red. Each bet has a payout designed so that the expected return is approximately \$0.947 per dollar wagered, resulting in an expected loss of \$-0.053 per play. This is modeled as an MDP with one state and $171$ actions, where one action is not placing a bet, ending the episode. If the player chooses not to bet, the payout is \$0. Assuming the player bets \$1 each time, the optimal strategy is to refrain from gambling, resulting in an expected profit of \$0. In this example, the choice of successive relaxation factor $w$ is known as we have only one state. More specifically, $0<w\leq w^{*}=\frac{1}{1-\gamma\min_{i,a}p^{i}_{ia}}=\frac{1}{1-\gamma}$. Therefore, both MF-DSORQL and MF-SORQL are the same as that of DSORQL and SORQL, respectively. The experiment is conducted with $\gamma=0.95$, and the polynomial step-size mentioned in [8]. Fig. 4,4,4, and 4, presents the behavior of the proposed algorithms for various choices of SOR parameters and compares them with QL and DQL. Each experiment consists of $100000$ trails, and the graph is plotted by taking the average over $10$ independent experiments. It is evident from the plot that the performance of DSORQL is consistent for various choices of $w$, whereas the performance of the SORQL algorithm improves as $w$ gets closer to $w^{*}$. Further, the over-estimation bias of the SORQL algorithm is visible when $w=1.3$ (Fig. 4) and $w=5$ (Fig. 4). It is also interesting to see that, in Fig. 4., the over-estimation bias is not visible as this is a special case wherein $w=w^{*}=\frac{1}{1-\gamma}$, and since the next state $j$ will be $i$, Equation (6) reduces to the following:
$Q_{n+1}(i,a)$ $=(1-\beta_{n}(i,a))Q_{n}(s,a)+\beta_{n}(i,a)\left(\frac{r^{i}_{ia}}{1-\gamma}\right).$
Therefore, in this particular case, the maximum operator gets canceled, and this is the reason for the behavior in Fig. 4. Note that in this classic example, the performance of the proposed DSORQL and DQL are almost the same.
V-A2 Grid World
Consider a $3\times 3$ grid world MDP. The agent can choose from four actions along four directions. The agent starts in the fixed starting state from one corner, with the goal state in the diagonally opposite corner. If the agent’s actions lead off the grid, it will stay in the same state. For every non-terminal step, the agent receives a random reward of either $-12$ or $+10$ with a probability of $0.5$ each. If the agent reaches the goal state, then any action from the goal state will end the episode with a reward of $+5$. The optimal policy concludes the episode after five actions, achieving an average reward of $+0.2$ per step. In order for this MDP to satisfy the requirement of $p^{i}_{ia}>0,\forall(i,a)\in S\times A$. The MDP is modified so that for any action chosen by the agent, there is a $0.9$ probability of transitioning to the next state according to the chosen action, and the agent stays in the same state with $0.1$ probability. Fig. 5 shows the average reward per step over $10000$ independent experiments. From Fig. 5, it is evident that the proposed MF-DSORQL outperforms MF-SORQL, QL, and DQL by obtaining a higher reward per step. Also, in this example, both QL and MF-SORQL show similar results, but neither is performing effectively.
Figure 5: The performance of algorithms on grid world example.
@@ V-B Deep RL version
This subsection compares the proposed double SOR Q-network (DSORDQN) with the deep RL version of SOR Q-learning (SORDQN) [17], the deep Q-network (DQN) [16], and the deep double Q-network (DDQN) [24]. The performance is evaluated using the maximization bias example [23], the LunarLander domain [22], and the classic CartPole environment [22]. The choice of SOR parameter $w$ for both SORDQN and DSORDQN is $1.3$ in all the experiments.
V-B1 Maximization Bias Example
@@
This example is similar to the classic scenario used to illustrate bias behavior in tabular QL and DQL, which is discussed in Chapter 6 of [2]. However, the current implementation is adapted from [23]. Specifically, the maximization bias example in this manuscript involves $\{0,1,2,...,M\}$ states with $M=10^{9}$ and one terminal state. Therefore, applying tabular methods is infeasible. Consequently, we apply the proposed DSORDQN algorithm and compare its performance with DQN, SORDQN, and DDQN. The agent always begins in state zero. The discount factor is $\gamma=0.999$. Each state allows two actions: moving left or right. If the agent moves right from state zero, it receives zero reward, and the game ends. If the agent moves left, it transitions to one of the states in $\{1,2,...,M\}$ with equal probability and receives zero reward. In these new states, moving right returns the agent to state zero, while moving left ends the game, with rewards drawn from a Gaussian distribution $N(-0.1,1)$. Note that the choice of parameters in this example is the same as that of [23]. In other words, a fully connected neural network with two hidden layers of dimension $4$ and $8$ is used to approximate the Q-function. The activation function is ReLU, and the optimizer is stochastic gradient descent with no momentum. The policy used is epsilon-greedy with $\epsilon=0.1$, with training conducted over $400$ episodes. We plot the probability of selecting the left action from state zero at the end of each episode, averaging over $1000$ runs. A higher probability of choosing left suggests that the algorithm is following a sub-optimal strategy, as consistently choosing right yields the highest mean reward. In Fig. 6, one can observe that the policy obtained from the proposed algorithm is better than DQN, SORDQN, and DDQN.
@@
V-B2 CartPole and LunarLander Environment
In the CartPole-v0 environment, the goal is to balance a pole on a moving cart by applying forces in the left and right direction on the cart. The agent receives a reward of $+1$ for each time step the pole remains upright, and the episode terminates if the pole angle is not in the range of ±$\pm 12^{\circ}$ or the cart range is not within $(-2.4,2.4)$. The standard threshold for a successful CartPole-v0 agent is to achieve an average reward of $195$ over $100$ episodes. In contrast, the LunarLander-v2 environment involves controlling a spacecraft to land safely on a lunar surface. The agent is rewarded for maintaining the lander within the landing pad, controlling its orientation, and landing with minimal velocity. The threshold for LunarLander-v2 is to achieve an average reward of $200$ over $100$ episodes. Both environments serve as tests for algorithms to demonstrate the ability to make decisions at varying complexity levels. The parameter selection and the implementation are taken from [26]. Specifically, the input layer has nodes of dimension state size, followed by two hidden layers with $64$ nodes each, using ReLU activation. The output layer has nodes of dimension action space. The optimizer used is Adam, and the policy is epsilon-greedy with decaying epsilon. The batch size is $64$, $\gamma$ is $0.99$, and the buffer size is $100,000$. Table I provides the performance of the proposed DSORDQN in comparison to the existing algorithms. It is evident from the table that the proposed algorithm completes both tasks with the minimum number of episodes compared to all the other algorithms.
Algorithm CartPole LunarLander
DQN 1000 1101
DDQN 3526 853
SORDQN 2343 826
DSORDQN 947 546
TABLE I: Number of episodes required to complete.
@@ VI Conclusion
Motivated by the recent success of SOR methods in Q-learning and the issues concerning its over-estimation bias, this manuscript proposes a double successive over-relaxation Q-learning. Both theoretically and empirically, the proposed algorithm demonstrates a lower bias compared to SOR Q-learning. In the tabular setting, convergence is proven under suitable assumptions using SA techniques. A deep RL version of the method is also presented. Experiments on standard examples, including roulette, grid-world, and popular gym environments, showcase the algorithm’s performance. It is worth mentioning that, similar to the boundedness of iterates observed in Q-learning [27], investigating the boundedness of the proposed algorithms in the tabular setting presents an interesting theoretical question for future research.

### v2

@@ IV-A Theoretical Analysis of Over-estimation Bias in SORQL
Bias $=\mathbb{E}[w\max_{i}\mu_{i}+(1-w)\max_{i}\mu_{i}]-\max_{i}\mathbb{E}[\mu_{i}]$ (36)
$=\mathbb{E}[\max_{i}\mu_{i}]-\max_{i}\mathbb{E}[\mu_{i}]\geq 0.$ (37)
Therefore, with probability $p^{i}_{ia}$, the SORQL algorithm overestimates the true value. Similarly, we can see that the MF-SORQL will also suffer from over-estimation bias. To this end, the techniques developed in double Q-learning prove useful. More precisely, in double Q-learning, for a set of random variables $\{X_{i}\}^{d}_{i=1}$, to evaluate $\max_{i}\mathbb{E}[X_{i}]$, two sets of unbiased estimators, namely $\{\mu^{A}_{i}\}^{d}_{i=1}$, and $\{\mu^{B}_{i}\}^{d}_{i=1}$ are used. Further, it is shown that for $a^{*}$ such that $\mu^{A}_{a^{*}}=\max_{i}\mu^{A}_{i}$, $\mathbb{E}[\mu^{B}_{a^{*}}]\leq\max_{i}\mathbb{E}[X_{i}]$ (Lemma 1, [11]). Therefore, Bias $=\mathbb{E}[\mu^{B}_{a^{*}}]-\max_{i}\mathbb{E}[\mu^{B}_{i}]\leq 0$, consequently, double Q-learning underestimates the true value. Note that underestimation is not a complete solution to the over-estimation problem of Q-learning. However, this modification to Q-learning has proven to be more useful in many situations [29]. Now, for the proposed DSORQL, let $\{\mu^{A}_{i}\}^{d}_{i=1}$, $\{\mu^{B}_{i}\}^{d}_{i=1}$, and $\{\zeta^{A}_{j}\}^{d}_{j=1}$, $\{\zeta^{B}_{j}\}^{d}_{j=1}$ be the unbiased estimators of $\{X_{i}\}^{d}_{i=1}$, and $\{Y_{j}\}^{d}_{j=1}$, respectively. Therefore, the bias is given by
$\text{Bias}=\mathbb{E}[w\mu^{B}_{a^{*}}+(1-w)\zeta^{B}_{b^{*}}]-\max_{i}\mathbb{E}[\mu^{B}_{i}]$ (38)
@@
IV-B Experimental Analysis of Over-estimation Bias in SORQL using Roulette [11]
In roulette, a player can choose from $170$ different betting actions, such as betting on a specific number or on the color black or red. Each bet has a payout designed so that the expected return is approximately \$0.947 per dollar wagered, resulting in an expected loss of \$-0.053 per play. This is modeled as an MDP with one state and $171$ actions, where one action is not placing a bet, ending the episode. If the player chooses not to bet, the payout is \$0. Assuming the player bets \$1 each time, the optimal strategy is to refrain from gambling, resulting in an expected profit of \$0. In this example, the choice of successive relaxation factor $w$ is known as we have only one state. More specifically, $0<w\leq w^{*}=\frac{1}{1-\gamma\min_{i,a}p^{i}_{ia}}=\frac{1}{1-\gamma}$. Therefore, both MF-DSORQL and MF-SORQL are the same as that of DSORQL and SORQL, respectively.
Figure 1: SORQL for different SOR parameters.
Figure 2: DSORQL for different SOR parameters.
The experiment is conducted with $\gamma=0.95$, and the polynomial step-size mentioned in [11]. Fig. 1 and Fig. 2 illustrate the performance of the SORQL and DSORQL algorithms, respectively, for different choices of SOR parameters, and compare them with QL [3] and DQL [11]. Each experiment involves $100000$ trials, and the graphs are obtained by averaging the results over $10$ independent runs. Following [11] all algorithms are updated synchronously and trained using epsilon-greedy policy. From Fig. 2, it is evident that the performance of DSORQL remains consistent with that of DQL for various values of $w$. In contrast, Fig. 1 clearly demonstrates the overestimation bias in SORQL algorithm.
@@ V Numerical Experiments
In this section, the tabular and deep RL versions of the proposed algorithms are compared on benchmark examples. The detailed experimental setup is in the supplementary material. The code for all the numerical experiments in this manuscript is publicly available and can be accessed in the author’s GitHub repository [30].
@@ V-A Tabular Version
This subsection provides numerical comparisons between Q-learning (QL) [3], double Q-learning (DQL) [11], triple-average policy gradient (TAPG) [13], model-free SOR Q-learning (MF-SORQL), and model-free double SOR Q-learning (MF-DSORQL). More specifically, we compare the above-mentioned algorithms on benchmark examples such as CartPole [18], LunarLander [18], and multi-armed bandit example discussed in [12], and [17]. The choice of SOR parameter for SORQL and DSORQL are kept same, and the parameter $\beta$ for TAPG is set to 0.95 as discussed in [13].
V-A1 CartPole Environment
In the CartPole-v0 environment, the goal is to balance a pole on a moving cart by applying forces in the left and right direction on the cart. The agent receives a reward of $+1$ for each time step the pole remains upright, and the episode terminates if the pole angle is not in the range of ±$\pm 12^{\circ}$ or the cart range is not within $(-2.4,2.4)$. In this experiment, the discount factor is 0.999, and the step size is $\frac{40}{n+100}$. We use an epsilon-greedy policy to train the algorithms. The continuous CartPole environment is discretized to $72$ states. The standard threshold for a successful CartPole-v0 agent is to achieve an average reward of $195$ over $50$ episodes. Table I, provides the minimum number of episodes required to achieve an average reward of $195$ over $50$ episodes. From Table I, it is evident that the proposed DSORQL algorithm, with the SOR parameter $w=1.1$, achieves the target much quicker than the rest of the algorithms.
V-A2 LunarLander Environment
The LunarLander-v2 environment involves controlling a spacecraft to land safely on a lunar surface. The agent is rewarded for maintaining the lander within the landing pad, controlling its orientation, and landing with minimal velocity. This environment serves as a test for algorithms to demonstrate the ability to make decisions at varying complexity levels. The threshold for LunarLander-v2 is to achieve a reward of $200$. The following table, i.e., Table I, provides the minimum number of episodes required to achieve the threshold averaged over five independent experiments. The step-size for all the algorithms are same with the discount factor is set to $0.99$. Once again, from Table I it is clear that the proposed algorithm with the SOR parameter $w=1.3$ learns the optimal policy much quicker than the rest of the algorithms.
Algorithm QL [3] SORQL [9] TAPG [13] DQL [11] DSORQL
CartPole 498 648 1000 1000 298
LunarLander 1065 1128 1073 1085 905
TABLE I: Number of episodes required to complete.
V-A3 Multi-Armed Bandit Example [12, 17]
This example considers a scenario with a single state and multiple action options. Specifically, there are thirty-nine available actions, i.e., $|A|=39$, and the thirty-ninth action leads to the end of the episode with no reward (See, Fig. 3). The other thirty-eight actions provide a reward sampled from a Gaussian distribution with a mean of $-0.0526$ and a variance of one. In this example, the discount factor, $\gamma$, is set to $0.99$, and the step size for all algorithms is defined as $100/(n+100)$. All state-action values are initialized to zero. Training is conducted over $50000$ episodes, averaging the results across $10$ independent runs. Similar to [12], all algorithms are updated asynchronously and trained using an exploratory policy.
The following table, i.e., Table II, presents the performance comparison of the proposed DSORQL algorithm with QL, DQL, SORQL, and TAPG. Note that the optimal Q-value in this example is zero.
Algorithm QL [3] SORQL [9] TAPG [13] DQL [11] DSORQL
$\max_{a}Q(a)$ 8.140 1.443 -5.27 -5.61 0.078
TABLE II: Approximate $Q^{*}(a)$ after 50,000 episodes.
From Table II, it is easy to see that DSORQL algorithm converges quicker to the optimal value in comparison to the rest of the algorithms. The choice of SOR parameter in this example for both SORQL and DSORQL is $w=1/(1-\gamma)$. Also, during our numerical experiments, we observed that this behavior of DSORQL remained consistent even when the cardinality of the action space was increased or decreased, as well as when the action distribution is modified to $N(-0.1,1)$, similar to [17]. This highlights the robustness of the proposed algorithm.
Figure 3: Multi-Armed Bandit Example [12].
@@ V-B Deep RL Version
This subsection compares the proposed double SOR Q-network (DSORDQN) with the deep RL version of SOR Q-learning (SORDQN) [16], the deep Q-network (DQN) [15], deep double Q-network (DDQN) [29], and Rainbow [31]. The performance is evaluated using the maximization bias example discussed in [19], and six Atari games. We have used MushroomRL library [32] to test the performance on the Atari games. The choice of SOR parameter $w$ in SORDQN and hyper-parameters in Rainbow are as discussed in [16] and [31], respectively.
V-B1 Maximization Bias Example [19]
@@
This example is similar to the classic scenario used to illustrate bias behavior in tabular QL and DQL, which is discussed in Chapter 6 of [2]. However, the current implementation is adapted from [19]. Specifically, the maximization bias example in this manuscript involves $\{0,1,2,...,M\}$ states with $M=10^{9}$ and two terminal state. Therefore, applying tabular methods is infeasible. Consequently, we apply the proposed DSORDQN algorithm and compare its performance with DQN, SORDQN, and DDQN. The agent always begins in state zero. The discount factor is $\gamma=0.999$. Each state allows two actions: moving left or right (See, Fig. 4). If the agent moves right from state zero, it receives zero reward, and the game ends. If the agent moves left, it transitions to one of the states in $\{1,2,...,M\}$ with equal probability and receives zero reward. In these new states, moving right returns the agent to state zero, while moving left ends the game, with rewards drawn from a Gaussian distribution $N(-0.1,1)$.
Note that the choice of parameters in this example is the same as that of [19] and discussed in detail in the supplementary material. The policy used is epsilon-greedy with $\epsilon=0.1$, with training conducted over $400$ episodes. We plot the probability of selecting the left action from state zero at the end of each episode, averaging over $1000$ runs. A higher probability of choosing left suggests that the algorithm is following a sub-optimal strategy, as consistently choosing right yields the highest mean reward. In Fig. 5, one can observe that the policy obtained from the proposed algorithm is better than DQN, SORDQN, and DDQN.
[…]
Figure 4: Any action from the states $1$ to $M$ gives reward sampled from $N(-0.1,1)$, where $M=10^{9}$.
V-B2 Atari Games
We evaluate the performance of DRL algorithms on six Atari 2600 games—Breakout, Qbert, Seaquest, Freeway, Asterix, and Space Invaders. The complete parameter set-up is available in the supplementary material. We report the mean reward over 10 epochs with standard deviation across five independent seeds, as summarized in Table III.
The results in Table III demonstrate that the proposed DSORDQN performs better compared to all baseline methods (DQN, SORDQN, DDQN) and Rainbow in four out of the five evaluated environments. Specifically, it achieves the highest average rewards in Seaquest (304.08 $\pm$ 5.64), Breakout (61.77 $\pm$ 2.75), Space Invaders (316.97 $\pm$ 15.01), Asterix (560.33 $\pm$ 15.03), and Freeway (25.04 $\pm$ 0.70). It also outperforms Rainbow by 28% in Seaquest, 219% in Breakout, 24% in Space Invaders, 7% in Asterix, and 67% in Freeway. In addition, it demonstrates lower variance across multiple runs, indicating more reliable learning outcomes.
While Rainbow achieves the highest reward in Qbert, it is a multi-step algorithm requiring more training samples, has several hyperparameters to tune, and our numerical experiments indicate that it takes longer to train due to its increased complexity. In contrast, DSORDQN provides a more efficient learning framework, yielding consistently better results across rest of the environments while maintaining computational feasibility.
@@ V-B2 Atari Games
We conclude this manuscript with a discussion on the ablation studies.
V-C Ablation Studies
In Fig. 1 and Fig. 2, we present a comparison between the SORQL and DSORQL algorithms across different values of the SOR parameter ($w$) on the roulette example. Specifically, we test $w=1.3$, $w=5$, and $w=\frac{1}{1-\gamma}$. The plots demonstrate that, unlike SORQL (Fig. 1), DSORQL consistently performs well across these varying choices of $w$ (Fig. 2). This result suggests that DSORQL is more robust to changes in $w$ compared to SORQL.
In the CartPole environment, we found that $w=1.1$ outperforms other choices of $w$ (e.g., $w=1.2$, $w=1.3$, and $w=1/(1-\gamma)$), which shows the importance of selecting the right $w$ for optimal performance. A similar trend is observed in the LunarLander domain, where $w=1.3$ delivers the best performance. For the multi-armed bandit task, we observe that performance improves as $w$ approaches $1/(1-\gamma)$.
In the maximization bias example, similar to the multi-armed bandit task, we again see performance improvements as $w$ is varied from $1.1$ to $1/(1-\gamma)$. For Atari games, experiments suggest that $w=1.2$ and $w=1.3$ yield promising performance on the discussed Atari tasks. Note that $w=1$ corresponds to DQL in the tabular setting and DDQN in the function approximation setting. From theory, it can be observed that faster convergence is achievable for $w>1$ [16], and hence, we choose the SOR parameter accordingly to optimize convergence. In our numerical experiments, we obtain a suitable SOR parameter by testing various values of $w$. Alternatively, hyperparameter tuning methods, such as those described in [33], can be applied to automate and refine this process.
@@ VI Conclusion
Motivated by the recent success of SOR methods in Q-learning and the issues concerning its over-estimation bias, this manuscript proposes a double successive over-relaxation Q-learning. Both theoretically and empirically, the proposed algorithm demonstrates a lower bias compared to SOR Q-learning. In the tabular setting, convergence is proven under suitable assumptions using SA techniques. A deep RL version of the method is also presented. Experiments on standard examples, showcase the algorithm’s performance. It is worth mentioning that, similar to the boundedness of iterates observed in Q-learning [34], investigating the boundedness of the proposed algorithms in the tabular setting presents an interesting theoretical question for future research.
Atari Games DQN [15] SORDQN [16] Rainbow [31] DDQN [29] DSORDQN
Seaquest 143.13 (37.85) 202.18 (19.77) 237.05 (12.20) 208.30 (60.02) 304.08 (5.64)
Breakout 55.34 (3.70) 4.27 (2.43) 19.36 (1.10) 54.02 (4.60) 61.77 (2.75)
Space Invaders 249.58 (11.62) 251.71 (14.69) 254.72 (8.70) 247.99 (31.03) 316.97 (15.01)
Qbert 552.78 (40.38) 267.18 (39.23) 2957.80 (86.33) 668.86 (58.26) 830.62 (31.71)
Asterix 437.59 (54.30) 236.62 (27.33) 524.51 (24.21) 408.11 (36.09) 560.33 (15.03)
Freeway 11.91 (3.88) 18.13 (1.62) 14.92 (1.79) 18.81 (1.06) 25.04 (0.70)
TABLE III: Mean test scores (with standard deviation in parentheses) over ten epochs, averaged across five random seeds.
Acknowledgement
I would like to sincerely thank Dr. Neha Bhadala and the anonymous referees for carefully reading the manuscript and providing constructive feedback, which greatly enhanced its presentation and overall quality.
Supplementary Material
This section provides details on the network architecture, training setup, and evaluation methodology for the deep reinforcement learning (DRL) experiments. Finally, the ablation studies for all numerical experiments are discussed.
VI-A Network Architecture, Training Setup, and Evaluation Methodology
Table I presents parameters for the maximization bias example, Table II presents the network architecture and training parameters for Atari, and Table III provides the parameters for the Rainbow, SORDQN, and DSORDQN algorithms.
TABLE IV: Neural network and training parameters used for the maximization bias example. The parameters are similar to [19].
Parameter Value
Network architecture Fully connected
Hidden layers 2
Layer sizes 4, 8
Activation ReLU
Optimizer SGD
Discount factor ($\gamma$) 0.999
Exploration Epsilon-greedy
Epsilon ($\epsilon$) 0.1
States $210^{9}+$ 2
Actions 2 (Left, Right)
Episodes 400
Iterations 1000
TABLE V: Network architecture and training parameters for Atari experiments.
Component Value
Env. 6 Atari games via MushroomRL
Input 84×84 grayscale (4-frame stack)
CNN 3 conv layers + 2 FC layers
Conv layers 32 (8×8/4), 64 (4×4/2), 64 (3×3/1)
FC layer 512 units, action-value output
Loss Huber (smooth L1)
Policy Epsilon-greedy ($\epsilon$: 1.0 → 0.1)
Eval $\epsilon$ 0.05
Hist. length 4
Train freq. 4 steps
Eval freq. 250,000 steps
Target update Every 10,000 steps
Replay init size 50,000
Max replay size 500,000
Test samples 125,000
Max train steps 2,500,000
Batch size 32
Discount factor ($\gamma$) 0.99
Optimizer Adam ($\alpha$ = 0.00025)
Metrics Mean reward (10 epochs), std over 5 seeds
TABLE VI: Parameters for Rainbow [31], SORDQN [16], and DSORDQN on Atari.
Parameter Value
Atoms ($n_{\text{atoms}}$) 51
Value range ($v_{\min}$, $v_{\max}$) (-10, 10)
Multi-step return ($n$) 3
Prioritization exponent ($\alpha$) 0.6
Priority correction ($\beta$) Linear (0.4 to 1)
Noisy net parameter ($\sigma$) 0.5
SOR param ($w$) 1.3